Whether or not a person has cheek dimples is an autosomally inherited character due to variation in a single gene. Having cheek dimples (D) is dominant to not having cheek dimple s(d). In a population of 28,000 Greek children, 3,500 had cheek dimples and the rest did not. Assuming that the population is in Hardy-Weingberg equilibrium,

(A) What are the frequencies of the two allels?

(B) How many of the children do you expect to be heterozygous?
Answer #1

As mentioned in the figure frequency of D is 1/8

so frequency of d is 1-1/8=7/8

so the frequency of Dd is=2*1/8*7/8=7/32

children heving heterozygous genotype is 28000*7/32=6125

so 6125 children will have heterozygous genotype.